Research Institute for Applied Knowledge-Processing (FAW)

Ulm Germany

Researchers Rüdiger Wirth.

ILP Areas Predicate invention, Applications of ILP.

Description of Research Research activities of FAW, mostly funded by research contracts from companies, cover a wide spectrum of applications of knowledge-based techniques to real world problems. The promotion of technology transfer from universities to companies is one of the most important tasks of FAW. One of the main problems addressed by most of the projects at FAW is how to provide knowledge in a form adequate for the problem at hand. Learning in various forms plays an important role in this task. In the context of ILP research FAW is investigating the integration of ILP techniques in knowledge-based systems solving to real world problems (e.g. [1]).

Rüdiger Wirth also contributes to the ESPRIT Basic Research Project 6020 “Inductive Logic Programming” as a consultant of the University of Stuttgart. His recent research activities include predicate invention [5], learning from positive examples only [3], and learning disjunctive definitions [2, 4].

Recent Publications

[1] M. Koegel. Aufbau einer Abstraktions-Hierarchie für strukturelle und funktionale Beschreibungen technischer Systeme (Constructing an abstraction hierarchy for structural and functional descriptions of technical devices), Diploma thesis, 1993 (in German).

[2] I. Stahl, B. Tausend, R. Wirth. Learning of disjunctive concepts via partitioning of the example set. Proc. Workshop on Logical Approaches to Machine Learning, European Conference on Artificial Intelligence, Vienna, 1992.

[3] I. Stahl, B. Tausend, R. Wirth. General-to-specific learning of Horn clauses from positive examples only. Proc. of IEEE 92 International Conference on Computer Systems and Software Engineering. the Hague, the Netherlands, 1992.

[4] I. Stahl, B. Tausend, R. Wirth. Two methods for improving inductive logic programming systems. Proc. European Conference on Machine Learning ECML-93, Vienna, Lecture Notes in Artificial Intelligence, Vol. 667. Springer-Verlag, 1993.

[5] R. Wirth, P. O’Rorke. Constraints for predicate invention. In S. Muggleton (ed.) Inductive Logic Programming. Academic Press, 1992.

Contact Person Rüdiger Wirth FAW Ulm P.O. Box 2060 D-89010 Ulm Germany Phone: +49 731 501 468 Fax: +49 731 501 999 E-mail: wirth@uni-ulm.de